A 3.6 μC charge is moved from a position where its electric potential energy is 7.2 mJ to a position where its potential energy is 1.8 mJ. What is the potential at the first position?
1
Expert's answer
2020-11-10T07:04:16-0500

The potential energy of an electrostatic field is

"E_{p,1} = q\cdot \varphi_1," where "\varphi" is the potential at the position in question. Therefore,

"\varphi_1 = \dfrac{E_{p,1}}{q} = \dfrac{7.2\cdot10^{-3}\,\mathrm{J}}{3.6\cdot10^{-6}\,\mathrm{C}} = 2\cdot10^3\,\mathrm{V}."

Worth noting that the potential is defined up to a constant, but it is often assumed that the potential is 0 when the distance tends to infinity.